Please help me solve this question thanks
Problem 3. The random variable X has density function f given by 0, elsewhere (a) Assuming that θ 0.8, determine K (b) Find Fx(t), the c.d.f. of X (c) Calculate P(0.4SX 0.8)
Answer #1

a)

for above to be valid: \int_{0}^{1} f(y) dy =1

\int_{0}^{1} f(y) dy =\int_{0}^{0.8}0.8y2 dy +\int_{0.8}^{1} Ky2 dy =0.8y3/3 |0.80 +Ky3/3 |10.8 =0.84/3 +K(1/3-0.83/3) =1

K=5.3082

b)

for 0 < y< 0.8

F(y) =\int_{0}^{y}0.8y2 dy =0.8y3/3 |y0 =0.8y3/3

for 0.8 <y<1

F(Y) =\int_{0}^{0.8}0.8y2 dy +\int_{0.8}^{y} 5.3082 y2 dy = 0.8y3/3 |0.80 +5.3082*y3/3 |y0.8 =5.3082*y3/3-0.7694

c)

P(0.4<Y<0.8)=F(0.8)-F(0.4)=0.8*0.83/3 -0.8*(0.4)3/3 =0.1195